# Counterfactual Credit Guided Bayesian Optimization

### v1

@@ 5 Theoretical Analysis
Assumptions 1 There exists $\beta_{t}>0$ such that the GP-UCB confidence event $\mathcal{E}_{t}:=\Big\{\,\lvert f(\mathbf{x})-\mu_{t}(\mathbf{x})\rvert\leq\beta_{t}\,\sigma_{t}(\mathbf{x})\ \ \forall\mathbf{x}\in\mathcal{X}\,\Big\}$ holds with probability at least $1-\delta$ [Srinivas et al., 2010]. And there exist constants $c_{0},c_{1}>0$ such that for the zero-mean posterior process $\varepsilon_{t}:=f_{t}-\mu_{t}$ we have $\mathbb{E}\!\big[\sup_{\mathbf{x}\in\mathcal{X}}\varepsilon_{t}(\mathbf{x})\,\bigm|\,\mathcal{D}_{t}\big]\leq c_{0}S_{t}$ [Adler and Taylor, 2007], and, conditionally on $\mathcal{D}_{t}$, the centered variables $Z_{t}^{(j)}-\mathbb{E}[Z_{t}^{(j)}\mid\mathcal{D}_{t}]$ are $(c_{1}S_{t})$-sub-Gaussian ($\psi_{2}$-norms are $\leq c_{1}S_{t}$[Ledoux and Talagrand, 2013].
@@
We then analyze the cumulative regret of the Credit-Weighted UCB variant of CCGBO. Our strategy closely follows the classic GP‐UCB analysis of Srinivas et al. [Srinivas et al., 2010] and [Hvarfner et al., 2022b], with the key modification that at each iteration $t$ the standard UCB acquisition is rescaled by a factor $w_{t}(\mathbf{x})=\bigl[c(\mathbf{x})\bigr]^{\frac{\tau}{\,1+(t/M)}},$ where $c(\mathbf{x})$ is the normalized counterfactual credit and $\tau>0$ is the decay parameter. We show that this rescaling only incurs a constant multiplicative penalty in the cumulative regret bound.
Assumption 2 (Lipschitz on derivatives) There exist constants $a,b>0$ such that for every $j=1,\dots,p$,
$\Pr\!\Bigl(\sup_{\mathbf{x}\in\mathcal{X}}\Bigl|\tfrac{\partial f}{\partial x_{j}}(\mathbf{x})\Bigr|>L\Bigr)\;\leq\;a\,\exp\bigl(-L^{2}/b^{2}\bigr).$ (7)
Under this assumption and the usual choice of $\beta_{t}=O\bigl(\sqrt{\log(t^{2}\pi^{2}/6\delta)}\bigr)$, with probability at least $1-\delta$ we have simultaneously for all $t$ and $\mathbf{x}$: $|f(\mathbf{x})-\mu_{t-1}(\mathbf{x})|\;\leq\;\beta_{t}\,\sigma_{t-1}(\mathbf{x}).$
@@
where $R_{N}=O\bigl(\sqrt{N\,\gamma_{N}\,\beta_{N}}\bigr)$ is the standard GP-UCB bound from [Srinivas et al., 2010] and $\gamma_{N}$ is the maximum information gain. In particular, since $\max_{t}B_{t}/A_{t}=B_{1}/A_{1}<\infty$, CCGBO matches the vanilla GP-UCB rate up to a constant factor that tends to 1 when $\lambda\to 0$ (so $A_{t}=B_{t}=1$) or $c_{\max}/c_{\min}\to 1$.
@@ 6 Experiments
6.1 Experimental Setup
We include the following baselines: First, Standard GP‑UCB [Srinivas et al., 2010] and Random Search are the most natural baselines for BO. Then, we include Weighted Gaussian Process UCB (WGP) [Deng et al., 2022] and Time-Varying Gaussian Process UCB (RGP) [Bogunovic et al., 2016] to cover the non‐stationary scenario. By comparing with WGP and RGP, we show that CCGBO’s credit mechanism can also eliminate low-value observations during the optimization process. Next, OutlierBO [Martinez-Cantin et al., 2018] represents robust outlier handling. In contrast, CCGBO’s credit assignment naturally downweights low‐contribution points (including noise and outliers). Finally, PiBO [Hvarfner et al., 2022b] and ColaBO [Hvarfner et al., 2024b] embody user‐prior, requiring expert‐supplied beliefs about the optimum. By comparing with PiBO and COLABO, we highlight that our can achieve competitive results without any external prior. WGP, RGP, and OutlierBO are grouped as knowledge from previous experiments methods, whereas PiBO and ColaBO belong to Embedding Priors over the Function Optimum methods. We demonstrate the superiority of our algorithm by comparing it with standard, non-stationary, robust, and prior guidance methods. Note on PiBO and ColaBO, for a fair comparison, we do not elicit any external priors from users. Instead, we construct the user‐provided prior by taking the location of the best point in the initial set as the prior mean for both PiBO and ColaBO.
In all experiments, each function evaluation is corrupted by independent Gaussian noise $\varepsilon\sim\mathcal{N}(0,\sigma_{n}^{2})$ with $\sigma_{n}^{2}=0.01$, to simulate realistic measurements. The size of the initial design is set adaptively to $n_{\mathrm{init}}=\max\bigl(2\,d,\;10\bigr)\,$, where $d$ denotes the dimensionality of the problem. We employ a Matérn $5/2$ kernel for the GP, with hyperparameters learned via maximum likelihood estimation. The iteration budget is 100. The regret is averaged on 50 independent experiments. We set the “half‐life” parameter $M=20$, the Monte Carlo UCB proxy sample number $K=25$, and the number of nearest neighbors $H=5$. We employed eight benchmarks, five synthetic and three real-world problems. For clarity, all the tasks aim for function maximization. All experiments were conducted on a MacBook Pro with Apple M2 Pro (10-core CPU, 16 GB RAM).
Synthetic Test Problems. We evaluate five synthetic test functions: Langermann2 is a 2-dimensional multimodal function on $0,10]^{2}$; Hartmann6 is a six-dimensional function on $0,1]^{6}$; Griewank6 is a nonconvex six-dimensional landscape defined on $600,600]^{6}$; Levy8 is an 8-dimensional valley-shaped function on $10,10]^{8}$; and Rosenbrock10 is a 10-dimensional multimodal test function on $5,10]^{10}$. We put high-dim synthetic test problems in Supplementary Materials.
Real-World Test Problems. We further evaluated CCGBO on three real-world problems. (1) The NAS experiment models a breast cancer prediction problem from the UCI repository [Dua and Graff, 2017] of searching for optimal hyperparameters as a BO problem. Specifically, each query $(x_{t},y_{t})$ corresponds to a choice of ([32,128],[1e6,1.0],[1e6,1.0],[1,8][32,128],[1e-6,1.0],[1e-6,1.0],[1,8] for batch size, Learning Rate, Learning Rate decay, hidden dim, respectively), where $y_{t}$ is the test classification error. (2) The Robot4 [Kaelbling and Lozano-Pérez, 2017] problem simulates a 4-dimensional robot pushing task using the Box2D physics engine, where each query corresponds to robot initial position $\in[-5,5]^{2}$ , initial robot angle control $\in[0,2\pi]$and simulation steps $\in[1,30]$, with $y_{t}$ being the Euclidean distance between the pushed object’s final position and target. (3) The Portfolio problem [Bruni et al., 2016] represents a 5-dimensional financial portfolio optimization task, where each query corresponds to asset normalized weight allocations $\in[0,1]^{5}$.
6.2 Comparison of CCGBO Against Baselines
@@
Figures 2 summarize the results on eight benchmarks, show the mean and standard error of the simregret and cumregret on 8 benchmark problems. We also report the area under simple regret in Table 1 as a summary metric of the entire optimization trajectory, defined as $\mathrm{AUSR}=\frac{1}{T-1}\sum_{t=2}^{T}\frac{r_{t-1}+r_{t}}{2},$ where $r_{t}$ denotes the simple regret at iteration $t$. It measures the remaining gap between the best value found so far and the true optimum. We report the observations below:
@@
(2) Lower cumulative regret: CCGBO maintains an advantage over most baselines, while random search quickly accumulates regret and specialized schemes like WGP and RGP only partially close the gap.This confirming that CCGBO preserves the sublinear regret rate while delivering practical gains.
@@
Our algorithm is plug-and-play, it also achieves good results on other different aquisition functions and high-dimensional problems. We put the experiment on other acquisition functions (TS, EI, JES etc.), high-dim problems (25 to 1000 dimension) and ablation study on $M$, $K$, $H$ in Supplementary Materials. Due to space limitations, details is provided in Supplementary Materials.
@@ B.1 Consistency of the MC proxy of the optimum
Assumptions 1 There exists $\beta_{t}>0$ such that the GP-UCB confidence event $\mathcal{E}_{t}:=\Big\{\,\lvert f(\mathbf{x})-\mu_{t}(\mathbf{x})\rvert\leq\beta_{t}\,\sigma_{t}(\mathbf{x})\ \ \forall\mathbf{x}\in\mathcal{X}\,\Big\}$ holds with probability at least $1-\delta$ [Srinivas et al., 2010]. And there exist constants $c_{0},c_{1}>0$ such that for the zero-mean posterior process $\varepsilon_{t}:=f_{t}-\mu_{t}$ we have $\mathbb{E}\!\big[\sup_{\mathbf{x}\in\mathcal{X}}\varepsilon_{t}(\mathbf{x})\,\bigm|\,\mathcal{D}_{t}\big]\leq c_{0}S_{t}$ [Adler and Taylor, 2007], and, conditionally on $\mathcal{D}_{t}$, the centered variables $Z_{t}^{(j)}-\mathbb{E}[Z_{t}^{(j)}\mid\mathcal{D}_{t}]$ are $(c_{1}S_{t})$-sub-Gaussian (equivalently, their $\psi_{2}$-norms are $\leq c_{1}S_{t}$[Ledoux and Talagrand, 2013].
@@ B.2 Regret Bound
We denote the instantaneous regret at step $t$ by $r_{t}=f(\mathbf{x}^{*})-f(\mathbf{x}_{t}),$ where $\mathbf{x}^{*}=\arg\max_{\mathbf{x}}f(\mathbf{x})$, and the cumulative regret by $R_{N}=\sum_{t=1}^{N}r_{t}.$ We will analyze the Credit-Weighted UCB (CCGBO) acquisition $\alpha_{t}(\mathbf{x})=\bigl(1-\lambda+\lambda\,w_{t}(\mathbf{x})\bigr)\,\bigl[\mu_{t-1}(\mathbf{x})+\beta_{t}\,\sigma_{t-1}(\mathbf{x})\bigr],$ where $w_{t}(\mathbf{x})=\bigl[c(\mathbf{x})\bigr]^{\tfrac{\tau}{1+t/M}},$ and we further assume $0<c_{\min}\leq c(\mathbf{x})\leq c_{\max}<\infty~\forall\mathbf{x}\in\mathcal{X},$ $\lambda\in[0,1]$ is the credit‐influence parameter, and $\{\beta_{t}\}$ is chosen as in [Srinivas et al., 2010]. We make the following standard assumption on the GP sample paths:
Assumption 2 (Lipschitz on derivatives) There exist constants $a,b>0$ such that for every $j=1,\dots,p$,
$\Pr\!\Bigl(\sup_{\mathbf{x}\in\mathcal{X}}\Bigl|\tfrac{\partial f}{\partial x_{j}}(\mathbf{x})\Bigr|>L\Bigr)\;\leq\;a\,\exp\bigl(-L^{2}/b^{2}\bigr).$ (11)
Under this assumption and the usual choice of $\beta_{t}=O\bigl(\sqrt{\log(t^{2}\pi^{2}/6\delta)}\bigr)$, with probability at least $1-\delta$ we have simultaneously for all $t$ and $\mathbf{x}$: $|f(\mathbf{x})-\mu_{t-1}(\mathbf{x})|\;\leq\;\beta_{t}\,\sigma_{t-1}(\mathbf{x}).$
@@
where $R_{N}=O\bigl(\sqrt{N\,\gamma_{N}\,\beta_{N}}\bigr)$ is the standard GP-UCB bound from [Srinivas et al., 2010] and $\gamma_{N}$ is the maximum information gain. In particular, since $\max_{t}B_{t}/A_{t}=B_{1}/A_{1}<\infty$, CCGBO matches the vanilla GP-UCB rate up to a constant factor that tends to 1 when $\lambda\to 0$ (so $A_{t}=B_{t}=1$) or $c_{\max}/c_{\min}\to 1$.
@@
Condition on the event $\mathcal{E}$ that for all $t,\mathbf{x}$, $f(\mathbf{x})\leq\mu_{t-1}(\mathbf{x})+\beta_{t}\,\sigma_{t-1}(\mathbf{x}),~f(\mathbf{x})\geq\mu_{t-1}(\mathbf{x})-\beta_{t}\,\sigma_{t-1}(\mathbf{x}).$ By definition, $\mathbf{x}_{t}=\arg\max_{\mathbf{x}}a_{t}(\mathbf{x})$, so for the true maximizer $\mathbf{x}^{*}$, $a_{t}(\mathbf{x}^{*})\leq a_{t}(\mathbf{x}_{t}).$ Thus
$\begin{split}&A_{t}\bigl[\mu_{t-1}(\mathbf{x}^{*})+\beta_{t}\,\sigma_{t-1}(\mathbf{x}^{*})\bigr]\\ &\leq\bigl(1-\lambda+\lambda\,w_{t}(\mathbf{x}^{*})\bigr)\bigl[\mu_{t-1}(\mathbf{x}^{*})+\beta_{t}\,\sigma_{t-1}(\mathbf{x}^{*})\bigr]\\ &\leq a_{t}(\mathbf{x}_{t})\leq B_{t}\bigl[\mu_{t-1}(\mathbf{x}_{t})+\beta_{t}\,\sigma_{t-1}(\mathbf{x}_{t})\bigr].\end{split}$ (13)
Rearranging gives $\mu_{t-1}(\mathbf{x}^{*})+\beta_{t}\,\sigma_{t-1}(\mathbf{x}^{*})\leq\Bigl(\tfrac{B_{t}}{A_{t}}\Bigr)\bigl[\mu_{t-1}(\mathbf{x}_{t})+\beta_{t}\,\sigma_{t-1}(\mathbf{x}_{t})\bigr].$ Since $f(\mathbf{x}_{t})\geq\mu_{t-1}(\mathbf{x}_{t})-\beta_{t}\,\sigma_{t-1}(\mathbf{x}_{t})$, we obtain

### v2

@@ 5 Theoretical Analysis
Assumptions 1 There exists $\beta_{t}>0$ such that the GP-UCB confidence event $\mathcal{E}_{t}:=\Big\{\,\lvert f(\mathbf{x})-\mu_{t}(\mathbf{x})\rvert\leq\beta_{t}\,\sigma_{t}(\mathbf{x})\ \ \forall\mathbf{x}\in\mathcal{X}\,\Big\}$ holds with probability at least $1-\delta$ (Srinivas et al., 2010). And there exist constants $c_{0},c_{1}>0$ such that for the zero-mean posterior process $\varepsilon_{t}:=f_{t}-\mu_{t}$ we have $\mathbb{E}\!\big[\sup_{\mathbf{x}\in\mathcal{X}}\varepsilon_{t}(\mathbf{x})\,\bigm|\,\mathcal{D}_{t}\big]\leq c_{0}S_{t}$ (Adler and Taylor, 2007), and, conditionally on $\mathcal{D}_{t}$, the centered variables $Z_{t}^{(j)}-\mathbb{E}[Z_{t}^{(j)}\mid\mathcal{D}_{t}]$ are $(c_{1}S_{t})$-sub-Gaussian ($\psi_{2}$-norms are $\leq c_{1}S_{t}$(Ledoux and Talagrand, 2013).
@@
We then analyze the cumulative regret of the Credit-Weighted UCB variant of CCGBO. Our strategy closely follows the classic GP‐UCB analysis of Srinivas et al. (Srinivas et al., 2010) and (Hvarfner et al., 2022b), with the key modification that at each iteration $t$ the standard UCB acquisition is rescaled by a factor $w_{t}(\mathbf{x})=\bigl[c(\mathbf{x})\bigr]^{\frac{\tau}{\,1+(t/M)}},$ where $c(\mathbf{x})$ is the normalized counterfactual credit and $\tau>0$ is the decay parameter. We show that this rescaling only incurs a constant multiplicative penalty in the cumulative regret bound.
Assumption 2 (RKHS regularity) The objective function $f$ lies in the reproducing kernel Hilbert space (RKHS) $\mathcal{H}_{k}$ associated with kernel $k$, with $\|f\|_{\mathcal{H}_{k}}\leq B$ and $k(\mathbf{x},\mathbf{x})\leq 1$ for all $\mathbf{x}\in\mathcal{X}$. By the reproducing property, this implies the deterministic bound $|f(\mathbf{x})|\leq B$ for all $\mathbf{x}\in\mathcal{X}$.
Under this assumption, following Theorem 3 of Srinivas et al. (2010), we choose $\beta_{t}$ such that $|f(\mathbf{x})-\mu_{t-1}(\mathbf{x})|\;\leq\;\beta_{t}\,\sigma_{t-1}(\mathbf{x})$ holds with probability at least $1-\delta$ over the observation noise.
We define the cumulative regret of CCGBO as $R_{N}^{\mathrm{ccg}}=\sum_{t=1}^{N}r_{t}^{\mathrm{ccg}},\quad\text{where }r_{t}^{\mathrm{ccg}}=f(\mathbf{x}^{*})-f(\mathbf{x}_{t})$ is the instantaneous regret at iteration $t$.
@@
where $R_{N}=O\bigl(\sqrt{N\,\gamma_{N}\,\beta_{N}}\bigr)$ is the standard GP-UCB bound from (Srinivas et al., 2010) and $\gamma_{N}$ is the maximum information gain. In particular, since $\max_{t}B_{t}/A_{t}=B_{1}/A_{1}<\infty$, CCGBO matches the vanilla GP-UCB rate up to a constant factor that tends to 1 when $\lambda\to 0$ (so $A_{t}=B_{t}=1$) or $c_{\max}/c_{\min}\to 1$.
Remark 2 (Practical Interpretation of the Constant Factor).
The constant factor $\max_{1\leq t\leq N}(B_{t}/A_{t})$ in Theorem 5.2 arises from a conservative worst-case analysis. In practice, this factor is benign for two reasons: (1) The ratio $B_{t}/A_{t}$ is monotonically decreasing in $t$ and converges to 1 as $t\to\infty$, since the exponent $\tau/(1+t/M)\to 0$ causes both $c_{\min}^{\tau/(1+t/M)}$ and $c_{\max}^{\tau/(1+t/M)}$ to approach 1. (2) The worst-case factor $B_{1}/A_{1}$ occurs only at $t=1$ which is a modest overhead. Thus, while the theoretical bound includes this constant, the credit weighting effectively degenerates in later iterations, allowing CCGBO to inherit the standard GP-UCB convergence behavior asymptotically.
@@ 6 Experiments
6.1 Experimental Setup
We include the following baselines: First, Standard GP‑UCB (Srinivas et al., 2010) and Random Search are the most natural baselines for BO. Then, we include Weighted Gaussian Process UCB (WGP) (Deng et al., 2022) and Time-Varying Gaussian Process UCB (RGP) (Bogunovic et al., 2016) to cover the non‐stationary scenario. By comparing with WGP and RGP, we show that CCGBO’s credit mechanism can also eliminate low-value observations during the optimization process. Next, OutlierBO (Martinez-Cantin et al., 2018) represents robust outlier handling. In contrast, CCGBO’s credit assignment naturally downweights low‐contribution points (including noise and outliers). Finally, PiBO (Hvarfner et al., 2022b) and ColaBO (Hvarfner et al., 2024b) embody user‐prior, requiring expert‐supplied beliefs about the optimum. By comparing with PiBO and COLABO, we highlight that our approach can achieve competitive results without any external prior. WGP, RGP, and OutlierBO are grouped as knowledge from previous experiments methods, whereas PiBO and ColaBO belong to Embedding Priors over the Function Optimum methods. We demonstrate the superiority of our algorithm by comparing it with standard, non-stationary, robust, and prior guidance methods. Note on PiBO and ColaBO, for a fair comparison, we do not elicit any external priors from users. Instead, we construct the user‐provided prior by taking the location of the best point in the initial set as the prior mean for both PiBO and ColaBO.
In all experiments, each function evaluation is corrupted by independent Gaussian noise $\varepsilon\sim\mathcal{N}(0,\sigma_{n}^{2})$ with $\sigma_{n}^{2}=0.01$, to simulate realistic measurements. The size of the initial design is set adaptively to $n_{\mathrm{init}}=\max\bigl(2\,d,\;10\bigr)\,$, where $d$ denotes the dimensionality of the problem. We employ a Matérn $5/2$ kernel for the GP, with hyperparameters learned via maximum likelihood estimation. The iteration budget is 100. The regret is averaged on 50 independent experiments. For the acquisition function, we set the UCB exploration parameter $\beta_{t}=2.576$, corresponding to the $99\%$ quantile of a standard normal distribution, and the credit influence parameter $\lambda=0.5$. We set the “half‐life” parameter $M=20$, the Monte Carlo UCB proxy sample number $K=25$, and the number of nearest neighbors $H=5$. We employed eight benchmarks, five synthetic and three real-world problems. For clarity, all the tasks aim for function maximization. All experiments were conducted on a MacBook Pro with Apple M2 Pro (10-core CPU, 16 GB RAM). Our code is publicly available in the repository: https://github.com/Qiyu-Wei/CCGBO.
Synthetic Test Problems. We evaluate five synthetic test functions: Langermann2 is a 2-dimensional multimodal function on $0,10]^{2}$; Hartmann6 is a six-dimensional function on $0,1]^{6}$; Griewank6 is a nonconvex six-dimensional landscape defined on $600,600]^{6}$; Levy8 is an 8-dimensional valley-shaped function on $10,10]^{8}$; and Rosenbrock10 is a 10-dimensional test function on $5,10]^{10}$. We put high-dim synthetic test problems in Supplementary Materials.
Real-World Test Problems. We further evaluated CCGBO on three real-world problems. (1) The NAS experiment models a breast cancer prediction problem from the UCI repository (Dua and Graff, 2017) of searching for optimal hyperparameters as a BO problem. Specifically, each query $(x_{t},y_{t})$ corresponds to a choice of ([32,128],[1e6,1.0],[1e6,1.0],[1,8][32,128],[1e-6,1.0],[1e-6,1.0],[1,8] for batch size, Learning Rate, Learning Rate decay, hidden dim, respectively), where $y_{t}$ is the test classification error. (2) The Robot4 (Kaelbling and Lozano-Pérez, 2017) problem simulates a 4-dimensional robot pushing task using the Box2D physics engine, where each query corresponds to robot initial position $\in[-5,5]^{2}$ , initial robot angle control $\in[0,2\pi]$ and simulation steps $\in[1,30]$, with $y_{t}$ being the Euclidean distance between the pushed object’s final position and target. (3) The Portfolio problem (Bruni et al., 2016) represents a 5-dimensional financial portfolio optimization task, where each query corresponds to asset normalized weight allocations $\in[0,1]^{5}$.
6.2 Comparison of CCGBO Against Baselines
Figures 2 summarize the results on eight benchmarks, and show the mean and standard error of the simregret and cumregret on 8 benchmark problems. We also report the area under simple regret in Table 1 as a summary metric of the entire optimization trajectory, defined as $\mathrm{AUSR}=\frac{1}{T-1}\sum_{t=2}^{T}\frac{r_{t-1}+r_{t}}{2},$ where $r_{t}$ denotes the simple regret at iteration $t$. It measures the remaining gap between the best value found so far and the true optimum. We report the observations below:
@@ 6.2 Comparison of CCGBO Against Baselines
(2) Lower cumulative regret: CCGBO maintains an advantage over most baselines, while random search quickly accumulates regret and specialized schemes like WGP and RGP only partially close the gap. This confirms that CCGBO preserves the sublinear regret rate while delivering practical gains.
@@
Our algorithm is plug-and-play, it also achieves good results on other different acquisition functions and high-dimensional problems. We put the experiment on other acquisition functions (TS, EI, JES etc.), high-dim problems (25 to 1000 dimensions) and ablation study on $M$, $K$, $H$ in Supplementary Materials. Due to space limitations, details are provided in Supplementary Materials.
@@ B.1 Consistency of the MC proxy of the optimum
Assumptions 1 There exists $\beta_{t}>0$ such that the GP-UCB confidence event $\mathcal{E}_{t}:=\Big\{\,\lvert f(\mathbf{x})-\mu_{t}(\mathbf{x})\rvert\leq\beta_{t}\,\sigma_{t}(\mathbf{x})\ \ \forall\mathbf{x}\in\mathcal{X}\,\Big\}$ holds with probability at least $1-\delta$ (Srinivas et al., 2010). And there exist constants $c_{0},c_{1}>0$ such that for the zero-mean posterior process $\varepsilon_{t}:=f_{t}-\mu_{t}$ we have $\mathbb{E}\!\big[\sup_{\mathbf{x}\in\mathcal{X}}\varepsilon_{t}(\mathbf{x})\,\bigm|\,\mathcal{D}_{t}\big]\leq c_{0}S_{t}$ (Adler and Taylor, 2007), and, conditionally on $\mathcal{D}_{t}$, the centered variables $Z_{t}^{(j)}-\mathbb{E}[Z_{t}^{(j)}\mid\mathcal{D}_{t}]$ are $(c_{1}S_{t})$-sub-Gaussian (equivalently, their $\psi_{2}$-norms are $\leq c_{1}S_{t}$(Ledoux and Talagrand, 2013).
@@
Remark 3 (RKHS Setting).
Theorem 5.1 remains valid under the RKHS setting of Assumption 2. The decomposition into $T_{1}$ (Monte Carlo error) and $T_{2}$ (model bias) still applies: $T_{1}$ depends solely on the GP surrogate model, while the bound on $T_{2}$ follows from the confidence event which holds under RKHS assumptions per Theorem 3 of Srinivas et al. (2010).
@@ B.2 Regret Bound
We denote the instantaneous regret at step $t$ by $r_{t}=f(\mathbf{x}^{*})-f(\mathbf{x}_{t}),$ where $\mathbf{x}^{*}=\arg\max_{\mathbf{x}}f(\mathbf{x})$, and the cumulative regret by $R_{N}=\sum_{t=1}^{N}r_{t}.$ We will analyze the Credit-Weighted UCB (CCGBO) acquisition $\alpha_{t}(\mathbf{x})=\bigl(1-\lambda+\lambda\,w_{t}(\mathbf{x})\bigr)\,\bigl[\mu_{t-1}(\mathbf{x})+\beta_{t}\,\sigma_{t-1}(\mathbf{x})\bigr],$ where $w_{t}(\mathbf{x})=\bigl[c(\mathbf{x})\bigr]^{\tfrac{\tau}{1+t/M}},$ and we further assume $0<c_{\min}\leq c(\mathbf{x})\leq c_{\max}<\infty~\forall\mathbf{x}\in\mathcal{X},$ $\lambda\in[0,1]$ is the credit‐influence parameter, and $\{\beta_{t}\}$ is chosen as in Srinivas et al. (2010). Under Assumption 2, the confidence bound $|f(\mathbf{x})-\mu_{t-1}(\mathbf{x})|\leq\beta_{t}\,\sigma_{t-1}(\mathbf{x})$ holds with probability at least $1-\delta$.
@@
where $R_{N}=O\bigl(\sqrt{N\,\gamma_{N}\,\beta_{N}}\bigr)$ is the standard GP-UCB bound from (Srinivas et al., 2010) and $\gamma_{N}$ is the maximum information gain. In particular, since $\max_{t}B_{t}/A_{t}=B_{1}/A_{1}<\infty$, CCGBO matches the vanilla GP-UCB rate up to a constant factor that tends to 1 when $\lambda\to 0$ (so $A_{t}=B_{t}=1$) or $c_{\max}/c_{\min}\to 1$.
@@
Condition on the event $\mathcal{E}$ that for all $t,\mathbf{x}$, $f(\mathbf{x})\leq\mu_{t-1}(\mathbf{x})+\beta_{t}\,\sigma_{t-1}(\mathbf{x}),~f(\mathbf{x})\geq\mu_{t-1}(\mathbf{x})-\beta_{t}\,\sigma_{t-1}(\mathbf{x}).$ Define the shifted acquisition function. Let $B_{\min,t}:=\min_{\mathbf{x}\in\mathcal{X}}\bigl[\mu_{t-1}(\mathbf{x})+\beta_{t}\,\sigma_{t-1}(\mathbf{x})\bigr],$ and set $\tilde{\alpha}_{t}(\mathbf{x})=\mu_{t-1}(\mathbf{x})+\beta_{t}\,\sigma_{t-1}(\mathbf{x})-B_{\min,t}\geq 0$. The credit-weighted acquisition becomes $a_{t}(\mathbf{x})=\bigl(1-\lambda+\lambda\,w_{t}(\mathbf{x})\bigr)\,\tilde{\alpha}_{t}(\mathbf{x}).$ By definition, $\mathbf{x}_{t}=\arg\max_{\mathbf{x}}a_{t}(\mathbf{x})$, so for the true maximizer $\mathbf{x}^{*}$, $a_{t}(\mathbf{x}^{*})\leq a_{t}(\mathbf{x}_{t}).$ Thus
$\begin{split}&A_{t}\bigl[\mu_{t-1}(\mathbf{x}^{*})+\beta_{t}\,\sigma_{t-1}(\mathbf{x}^{*})-B_{\min,t}\bigr]\\ &\leq\bigl(1-\lambda+\lambda\,w_{t}(\mathbf{x}^{*})\bigr)\bigl[\mu_{t-1}(\mathbf{x}^{*})+\beta_{t}\,\sigma_{t-1}(\mathbf{x}^{*})-B_{\min,t}\bigr]\\ &=a_{t}(\mathbf{x}^{*})\leq a_{t}(\mathbf{x}_{t})\leq B_{t}\bigl[\mu_{t-1}(\mathbf{x}_{t})+\beta_{t}\,\sigma_{t-1}(\mathbf{x}_{t})-B_{\min,t}\bigr].\end{split}$ (11)
Rearranging gives $\mu_{t-1}(\mathbf{x}^{*})+\beta_{t}\,\sigma_{t-1}(\mathbf{x}^{*})-B_{\min,t}\leq\Bigl(\tfrac{B_{t}}{A_{t}}\Bigr)\bigl[\mu_{t-1}(\mathbf{x}_{t})+\beta_{t}\,\sigma_{t-1}(\mathbf{x}_{t})-B_{\min,t}\bigr].$ Since $f(\mathbf{x}^{*})\leq\mu_{t-1}(\mathbf{x}^{*})+\beta_{t}\,\sigma_{t-1}(\mathbf{x}^{*})$ and $f(\mathbf{x}_{t})\geq\mu_{t-1}(\mathbf{x}_{t})-\beta_{t}\,\sigma_{t-1}(\mathbf{x}_{t})$, we obtain